# A Benchmark for Multi-modal Foundation Models on Low-level Vision: from Single Images to Pairs

### v1

@@ III-B2 Description for image pairs
We also include the mPLUG-Owl2 (Q-Instruct) for description-pair ability benchmark. As shown in Table V, similarly, all MLLMs perform better in the aspect of relevance than completeness and precision. Furthermore, fine-tuned with the single image low-level visual dataset Q-Instruct [67], mPLUG-Owl2 (Q-Instruct) achieves the best performance on completeness and relevance but gets the lowest score on precision. This indicates the knowledge infusion from single images can effectively enhance an MLLM to focus on corresponding low-level dimensions for targeted responses, but it does not improve the accuracy of the content, meaning it cannot enhance the core analytical ability of image pairs.
@@ III-C Findings on Assessment
| Shikra (Vicuna-7B) [64] | argmax | 0.178/0.201 | 0.277/0.281 | 0.152/0.169 | 0.248/0.267 | 0.071/0.065 | 0.513/0.562 | 0.245/0.246 |
| Shikra (Vicuna-7B) [64] | softmax | 0.314/0.307 | 0.327/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 |
@@
| InstructBLIP (Vicuna-7B) [5] | softmax | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 |
@@
| mPLUG-Owl (LLaMA-7B) [52] | softmax | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 |
@@
| LLaMA-Adapter-V2 [66] | softmax | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.251/0.257 | 0.604/0.666 | 0.412/0.425 |
@@
| LLaVA-v1 (Vicuna-13B) [3] | softmax | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.285/0.297 | 0.626/0.684 | 0.349/0.372 |
@@ III-C1 MLLM Performance
To measure the assessment ability, we evaluate the performance of 20 open-source MLLMs on 7 IQA datasets that are with at least 1,000 images and 15 human ratings per image [70]. The experimental results are illustrated in Table VI. a) Primarily, we notice that the majority of MLLMs are notably better than NIQE on non-natural circumstances (CGI, AIGC, artificial distortions), showing their potential towards general-purpose evaluators on a broader range of low-level appearances. b) We also notice that without explicit alignment with human opinions during training, the most excellent MLLM, Emu2-Chat (which is based on the heaviest LLM, LLaMA-33B), can already outperform CLIP-ViT-Large-14 by a large margin (25%). These results have demonstrated that, though most MLLMs are still based on CLIP as visual encoders, their high capacity in the strong language decoder can do help them perform much better on visual quality assessment even without any explicit training.
@@ III-C2 Superiority of softmax
In this section, we quantitatively evaluate the correlation with human perception on a simple argmax strategy between goodbad and our proposed softmax strategy. In Table VII, we select 5 MLLMs of different architectures and confirm that for all IQA datasets, the more measurable softmax strategy predicts better than the argmax strategy, which degenerates into only two scores, 0 and 1. Though the result is generally expected, the experiments validate that MLLMs have quantitative assessment ability hidden behind their word outputs, and prove the effectiveness of our softmax-based IQA strategy.
@@
| InternLM-XComposer-VL (InternLM) [61] | ||||||||
@@ III-C3 Prompt Ensemble Effectiveness
As shown in Table VIII, the prompt ensemble strategy (as proposed in Eq. 2) on top-7 MLLMs (i.e. Emu2-Chat, InternLM-XComposer-VL, QWen-VL, InfiMM, LLaVA-v1.5 (13B), mPLUG-Owl, and LLaVA-v1.5 (7B)) can lead to up to 5% accuracy improvement (in average 1.7%). We believe it is a useful boost technique to improve the performance of MLLMs on the IQA task. Nevertheless, we also notice that different MLLMs perform best with different specific prompt combos. For example, the good+finepoor+bad performs best on InternLM-XComposer-VL, but comes with reduced accuracy on QWen-VL compared with only goodpoor. While goodpoor is proved overall best single word pair (except Emu2-Chat and LLaVA-v1.5 (13B)) for the evaluation and shows stable results across MLLMs, we decide to keep the current strategy (using the goodpoor combo) in Q-Bench.
@@ IV Conclusion
In this research, we introduce Q-Bench, a benchmark designed to evaluate the advancements of MLLMs in low-level visual abilities. We evaluate the MLLMs from three aspects: perception of low-level visual attributes, description of low-level visual content, and assessment of image quality. Additionally, acknowledging the importance of discerning differences and similarities in image pairs, our benchmark encompasses both single images and image pairs in the perception and description tasks. To evaluate these abilities, we have compiled two multi-modal benchmark datasets focused on low-level vision, and introduced a unified softmax-based method for quantitative image quality assessment (IQA) in MLLMs. Our findings demonstrate that some advanced MLLMs exhibit commendable low-level visual abilities even without specialized low-level training. However, there’s still a significant journey ahead before MLLMs can become fully reliable assistants in general low-level visual tasks. We hope that the insights gained from Q-Bench will spur further development in MLLMs, particularly in improving their perception and understanding of low-level visual elements.

### v2

@@ III-B2 Description for image pairs
We also include the mPLUG-Owl2 (Q-Instruct) for description-pair ability benchmark. As shown in Table V, similarly, all MLLMs perform better in the aspect of relevance than completeness and precision. Furthermore, fine-tuned with the single image low-level visual dataset Q-Instruct [q-instruct], mPLUG-Owl2 (Q-Instruct) achieves the best performance on completeness and relevance but gets the lowest score on precision. This indicates the knowledge infusion from single images can effectively enhance an MLLM to focus on corresponding low-level dimensions for targeted responses, but it does not improve the accuracy of the content, meaning it cannot enhance the core analytical ability of image pairs.
@@ III-C Findings on Assessment
| Shikra (Vicuna-7B) [shikra] | argmax | 0.178/0.201 | 0.277/0.281 | 0.152/0.169 | 0.248/0.267 | 0.071/0.065 | 0.513/0.562 | 0.245/0.246 |
| Shikra (Vicuna-7B) [shikra] | softmax | 0.314/0.307 | 0.327/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 |
@@
| InstructBLIP (Vicuna-7B) [iblip] | softmax | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 |
@@
| mPLUG-Owl (LLaMA-7B) [mplugowl] | softmax | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 |
@@
| LLaMA-Adapter-V2 [llamaadapterv2] | softmax | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.251/0.257 | 0.604/0.666 | 0.412/0.425 |
@@
| LLaVA-v1 (Vicuna-13B) [llava] | softmax | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.285/0.297 | 0.626/0.684 | 0.349/0.372 |
@@ III-C1 MLLM Performance
To measure the assessment ability, we evaluate the performance of 20 open-source MLLMs on 7 IQA datasets that are with at least 1,000 images and 15 human ratings per image [itu]. The experimental results are illustrated in Table VI. a) Primarily, we notice that the majority of MLLMs are notably better than NIQE on non-natural circumstances (CGI, AIGC, artificial distortions), showing their potential towards general-purpose evaluators on a broader range of low-level appearances. b) We also notice that without explicit alignment with human opinions during training, the most excellent MLLM, Emu2-Chat (which is based on the heaviest LLM, LLaMA-33B), can already outperform CLIP-ViT-Large-14 by a large margin (25%). These results have demonstrated that, though most MLLMs are still based on CLIP as visual encoders, their high capacity in the strong language decoder can do help them perform much better on visual quality assessment even without any explicit training.
@@ III-C2 Superiority of softmax
In this section, we quantitatively evaluate the correlation with human perception on a simple argmax strategy between goodbad and our proposed softmax strategy. In Table XII, we select 5 MLLMs of different architectures and confirm that for all IQA datasets, the more measurable softmax strategy predicts better than the argmax strategy, which degenerates into only two scores, 0 and 1. Though the result is generally expected, the experiments validate that MLLMs have quantitative assessment ability hidden behind their word outputs, and prove the effectiveness of our softmax-based IQA strategy.
@@
| InternLM-XComposer-VL (InternLM) [xcomposer] | ||||||||
@@ III-C3 Prompt Ensemble Effectiveness
As shown in Table XIII, the prompt ensemble strategy (as proposed in Eq. 2) on top-7 MLLMs (i.e. Emu2-Chat, InternLM-XComposer-VL, QWen-VL, InfiMM, LLaVA-v1.5 (13B), mPLUG-Owl, and LLaVA-v1.5 (7B)) can lead to up to 5% accuracy improvement (in average 1.7%). We believe it is a useful boost technique to improve the performance of MLLMs on the IQA task. Nevertheless, we also notice that different MLLMs perform best with different specific prompt combos. For example, the good+finepoor+bad performs best on InternLM-XComposer-VL, but comes with reduced accuracy on QWen-VL compared with only goodpoor. While goodpoor is proved overall best single word pair (except Emu2-Chat and LLaVA-v1.5 (13B)) for the evaluation and shows stable results across MLLMs, we decide to keep the current strategy (using the goodpoor combo) in Q-Bench+.
@@ IV Conclusion
In this research, we introduce Q-Bench+, a benchmark designed to evaluate the advancements of MLLMs in low-level visual abilities. We evaluate the MLLMs from three aspects: perception of low-level visual attributes, description of low-level visual content, and assessment of image quality. Additionally, acknowledging the importance of discerning differences and similarities in image pairs, our benchmark encompasses both single images and image pairs in the perception and description tasks. To evaluate these abilities, we have compiled two multi-modal benchmark datasets focused on low-level vision, and introduced a unified softmax-based method for quantitative image quality assessment (IQA) in MLLMs. Our findings demonstrate that some advanced MLLMs exhibit commendable low-level visual abilities even without specialized low-level training. However, there’s still a significant journey ahead before MLLMs can become fully reliable assistants in general low-level visual tasks. We hope that the insights gained from Q-Bench+ will spur further development in MLLMs, particularly in improving their perception and understanding of low-level visual elements.
a
V Appendix
V-A Subjective Experiment
A total of twenty experts, each with professional skills and extensive experience in photography, are invited to participate in the subjective labeling experiment of Q-Bench+. The subjective experiment takes place in a laboratory environment with standard indoor lighting. A Dell-4K monitor, which supports a resolution of , is used for displaying the interfaces. The screenshots of interfaces can be referred to in Fig. 6. Each expert annotates up to 30 images a day to avoid fatigue, and every annotation is carefully reviewed by at least three other experts before acceptance. In this way, we ensure the accuracy and rigor of the Q-Bench+ labels to the greatest extent possible. This, in turn, makes the performance testing capability of Q-Bench+ more precise and meaningful.
V-B Distortions and Other Low-level Attributes Enumeration
V-B1 Distortions
Blurs [lens blur (out-of-focus), motion blur, zoom blur, gaussian blur, glass blur], Noises [gaussian noise, speckle noise, pepper noise], Artifacts [compression artifact, transmission error], Exposure Issues [under-exposure, over-exposure], Miscellaneous Artificial Distortions [pixelate, color-diffusion, jitter, etc]
V-B2 Other low-level attributes
Color [color style, color vividity], Lighting [bright, dim], Composition [Symmetrical, Rule-of-Thirds], Visual Styles [animation, realism, computer-generated, AI-generated], Photographic Methods [background bokeh (shallow DOF), high contrast, motion blur (on fast-moving objects), etc]
V-C Multi-choice Question for Kosmos-2
While Kosmos-2 performs generally well on the description and assessment tasks, we notice that it is hardly capable of answering a multi-choice question with the general prompt form applicable for other methods, as follows:
How is the clarity of the image?
(Question) [IMAGE_TOKEN] (Image)
Choose between one of the following options:
A. High (Correct) B. Medium(Wrong) C. Low(Wrong)
For most situations (86%) in our primary sample test with the prompts above, Kosmos-2 will directly append a new candidate (e.g., D. Excellent or D. Very Low) answer instead of choosing one option among them, denoted as prompt failure. This might be because the language model of Kosmos-2 has smaller capacity (1B) than other MLLMs that are based on LLaMA/MPT (7B/13B).
Considering that the prompt failure is actually not directly related with low-level perception, we try different prompt engineering techniques to reduce the prompt failure rate, and finalize with a simple modification which can limit the prompt failure to less than 10% in our sample set, as follows:
How is the clarity of the image?
(Question) [IMAGE_TOKEN] (Image)
Choose between one of the following options:
A. High (Correct) B. Medium(Wrong) C. Low(Wrong)
#Answer:
Nevertheless, we are still not able to eliminate the prompt failures for Kosmos-2. Henceforth, to systematically remove the negative effect of prompt failures on multi-choice questions for Kosmos-2, we conduct a choice-free special setting for it, i.e. close-set inference, via ranking the perplexity of different answers and choose the answer with minimum generative loss:
How is the clarity of the image?
[IMAGE_TOKEN] #Answer: High
loss:7.43 ✓ Choose this.
How is the clarity of the image?
[IMAGE_TOKEN] #Answer: Medium
loss:7.56 ✗
How is the clarity of the image?
[IMAGE_TOKEN] #Answer: Low
loss:7.92 ✗
As shown in Table IX, perplexity-based close-set inference can notably improve results of Kosmos-2. Considering that it is still the MLLM with fewest parameters among the ten models, its results are decent at its model size. More importantly, they validate that our observation of the prompt failure is reasonable.
| Sub-categories | Question Types | Quadrants of Low-level Concerns | Overall | # | |||||
[…]
| Distortion | Other | ||||||||
| random guess | 50.00% | 28.18% | 33.30% | 37.54% | 38.49% | 38.70% | 36.50% | 37.87% | - |
[…]
| ⋆⋆Kosmos-2 (close-set) | 61.48% | 37.13% | 40.76% | 40.04% | 50.88% | 45.30% | 58.15% | 47.26% | ✓ |
| Model (variant) | all | correct answer: Yes | correct answer: No | mean | de-biased # |
| random guess | 50.00% | 50.00% | 50.00% | 50.00% | - |
[…]
| LLaVA-v1 (Vicuna-13B) | 57.10% | 60.29% | 51.66% | 55.97% | 6 |
| MiniGPT-4 (Vicuna-13B) | 57.56% | 70.00% | 37.38% | 53.69% | 8 |
| LLaMA-Adapter-V2 | 67.12% | 68.80% | 64.76% | 66.78% | 2 |
| InstructBLIP (Flan-T5-XL) | 68.67% | 80.14% | 50.23% | 65.19% | 4 |
| InstructBLIP (Vicuna-7B) | 71.40% | 84.32% | 50.47% | 67.39% | 1 |
| Otter-v1 (MPT-7B) | 57.74% | 70.14% | 37.38% | 53.76% | 7 |
| IDEFICS-Instruct (LLaMA-7B) | 59.74% | 88.65% | 13.09% | 50.87% | 9 |
| mPLUG-Owl (LLaMA-7B) | 69.31% | 95.82% | 26.67% | 61.25% | 5 |
V-D “Yes or No?”: How Biased are MLLMs?
In this section, we take a deeper analysis on the Yes-or-No judgment ability of MLLMs, that whether these models can get similar accuracy on questions that should be answered with Yes, as those should be replied as No. Sadly, we notice that all MLLMs have higher prediction accuracy on Yes-questions than No-questions, while some MLLMs are more very severe biased (e.g., IDEFICS-Instruct). Considering that our LLVisionQA dataset contains more (62%) Yes-questions than No-questions (38%) and may introduce biases while comparing different MLLMs, we further compute a de-biased accuracy for all these methods, as the mean value of the accuracies on two types of questions, and present the respective de-biased rank for all participating MLLMs, as listed in Table X. We hope this study on the biases and the de-biased results can provide a fairer comparison among them, as well as bring insights on the future improvements of MLLMs for low-level perception.
V-E Settings for GPT Evaluation of Perception
Given GPT’s inherent variability, identical prompts can yield non-definitive responses. To address the impact of such situations on our evaluation, we’ve implemented a 5-round voting strategy. Under this approach, we pose the same prompt as defined in the following templates five times, taking the popular votes of GPT’s answers to determine the final outcome. Our human analysis on a sample set confirms that the 5-round voting strategy improves GPT evaluation accuracy from 93.2% to 98.4%, reducing errors to only 1/4 compared with the single-round evaluation.
Prompt Templates for GPT Evaluation:
#System: You are a helpful assistant that grades answers related to image quality and aesthetics. There are a lot of special terms or keywords related to image processing and photography. You will pay attention to the context of ’quality evaluation’ when grading.
#User: Assuming you are a grader, you will now be provided with a question [question] and a set of options [options] with option [options[0]] being the correct answer. Additionally, there will be an answer [answer] provided by a respondent. Please determine whether the respondentś answer is correct considering the context of the question. Even if the word choice is not completely the same, you can decide based on the given options and see whether the one in the answer is close enough to the given correct answer, The result is 1 if the answer is correct and else the result is 0. Please only provide the result in the following format: Result:
V-F Comparing natural images and AI-generated images
There might be doubt that pairwise comparison between natural images (NIs) and AI-generated images (AIGIs) is unnecessary since the criteria for assessing quality differ between these two types of images is different. However, we believe it is essential to continue conducting comparisons between NIs and AIGIs for several reasons:
Common Technical Quality Issues: AIGIs are generated based on prior knowledge learned from NIs and, as such, they may inherit common technical issues such as noise, blur, and incorrect exposure. Therefore, examining and comparing the technical quality of NIs and AIGIs is meaningful, as it can guide and improve the technical quality of AIGIs.
Aesthetic Quality Comparability: AIGIs and NIs also share comparability in aesthetic quality, such as in composition, color adjustment, and style. Investigating the aesthetic quality comparison between AIGIs and NIs can enhance the corresponding capabilities of MLLMs, thus improving the aesthetic quality of AI-generated content. Furthermore, considering the greater aesthetic freedom and diversity of AIGIs, comparing them with NIs can also provide more comprehensive aesthetic guidance for the photography of NIs.
V-G Settings for GPT Evaluation of Description
The detailed input prompt template for GPT to evaluate the three dimensions is listed as follows:
#System: You are a helpful assistant.
Completeness. #User: Evaluate whether the description [MLLM_DESC] completely includes the low-level visual information in the reference description [GOLDEN_DESC].
Please rate score 2 for completely or almost completely including reference information, 0 for not including at all, 1 for including part of the information or similar description.
Please only provide the result in the following format: Score: -
Preciseness. #User Evaluate whether the description [MLLM_DESC] precisely reflects the reference description [GOLDEN_DESC].
Please rate score 2 for totally no controversial low-level description, 1 for less controversial low-level description than matched descrpition, and 0 for more controversial low-level description than matched description. Please only provide the result in the following format: Score: -
Relevance. #User: Evaluate whether the description [MLLM_DESC] is relevant to the low-level visual information, which may include blur, noise, exposure, artifact, color, lighting, focus, composition, etc.
Please rate score 2 for completely relevant, 1 for partly relevant, and 0 for totally irrelevant.
Please only provide the result in the following format: Score:
In the prompt template, the [MLLM_DESC] denotes the output description from MLLMs, and [GOLDEN_DESC] denotes the golden description in the LLDescribe+ dataset.
| Dataset Type | Prompt | In-the-wild | Generated | Artificial | Average | ||||
[…]
| InfiMM (Zephyr-7B) [InfiMM] | Prompt1 | 0.471/0.541 | 0.633/0.610 | 0.257/0.276 | 0.576/0.576 | 0.263/0.261 | 0.731/0.775 | 0.491/0.461 | 0.489/0.500 |
| Prompt2 | 0.496/0.533 | 0.605/0.618 | 0.266/0.291 | 0.569/0.593 | 0.239/0.248 | 0.708/0.768 | 0.492/0.490 | 0.482/0.506 | |
[…]
| Prompt2 | 0.694/0.712 | 0.732/0.738 | 0.363/0.366 | 0.644/0.613 | 0.321/0.342 | 0.779/0.772 | 0.844/0.820 | 0.625/0.623 | |
[…]
| Prompt2 | 0.571/0.621 | 0.728/0.748 | 0.360/0.410 | 0.629/0.683 | 0.236/0.261 | 0.734/0.773 | 0.521/0.538 | 0.540/0.576 | |
[…]
| Prompt2 | 0.512/0.513 | 0.443/0.465 | 0.303/0.315 | 0.408/0.415 | 0.318/0.324 | 0.697/0.752 | 0.392/0.421 | 0.439/0.458 | |
[…]
| Prompt2 | 0.474/0.498 | 0.565/0.588 | 0.314/0.345 | 0.488/0.521 | 0.311/0.322 | 0.692/0.771 | 0.382/0.392 | 0.461/0.491 | |
[…]
| Prompt2 | 0.541/0.600 | 0.632/0.617 | 0.286/0.316 | 0.570/0.577 | 0.301/0.318 | 0.664/0.719 | 0.416/0.429 | 0.487/0.511 | |
[…]
| Prompt2 | 0.395/0.421 | 0.633/0.647 | 0.233/0.258 | 0.455/0.496 | 0.147/0.173 | 0.685/0.704 | 0.463/0.483 | 0.430/0.455 | |
| Dataset Type | In-the-wild | Generated | Artificial | |||||
[…]
| Shikra (Vicuna-7B) [shikra] | argmax | 0.178/0.201 | 0.277/0.281 | 0.152/0.169 | 0.248/0.267 | 0.071/0.065 | 0.513/0.562 | 0.245/0.246 |
| Shikra (Vicuna-7B) [shikra] | softmax | 0.314/0.307 | 0.327/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 |
[…]
| InstructBLIP (Vicuna-7B) [iblip] | softmax | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 |
[…]
| mPLUG-Owl (LLaMA-7B) [mplugowl] | softmax | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 |
[…]
| LLaMA-Adapter-V2 [llamaadapterv2] | softmax | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.251/0.257 | 0.604/0.666 | 0.412/0.425 |
[…]
| LLaVA-v1 (Vicuna-13B) [llava] | softmax | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.285/0.297 | 0.626/0.684 | 0.349/0.372 |
| Dataset Type | In-the-wild | Generated | Artificial | Average | ||||
[…]
| InfiMM (Zephyr-7B) [InfiMM] | ||||||||
[…]
| Emu2-Chat (LLaMA-33B) [emu2] | ||||||||
[…]
| InternLM-XComposer-VL (InternLM) [xcomposer] | ||||||||
[…]
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | ||||||||
[…]
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | ||||||||
[…]
| Qwen-VL (QwenLM) [Qwen-VL] | ||||||||
[…]
| mPLUG-Owl (LLaMA-7B) [mplugowl] | ||||||||
[…]
V-H Evaluation Details for Assessment Ability
V-H1 Example pseudo code for MLLMs on IQA
In Algo. 1, we provide an example of how to evaluate image quality with MLLMs. The algorithm is simple with only 9 lines, and could be easily integrated with any new MLLMs (based on causal LLMs), so as to allow these models to quantitatively predict the quality of images.
V-H2 Detailed performance for softmax & prompt ensemble strategy
In this section, we present a detailed analysis of the performance of the softmax and prompt ensemble strategies. This comprehensive evaluation aims to provide a precise comparison and demonstrate the effectiveness of the proposed strategies. The softmax and argmax comparison is illustrated in Table XII while the prompt ensemble performance is exhibited in Table XIII.
V-H3 Enriching prompt set
We concur that incorporating more complex synonyms in the prompt ensemble strategy can help mitigate word bias and enable the prompt-ensemble strategy to more accurately target positive or negative quality embeddings. At the same time, we remain open to exploring different choices for prompt ensembles. The selection of terms in our paper represents a preliminary attempt, thus our choices are intentionally simplistic, aimed only at demonstrating the effectiveness of the prompt-ensemble strategy. Further enriching this selection is a valuable direction in prompt engineering, potentially unlocking the full assessment potential of the model. Therefore, we follow your suggestions by enriching the prompts with more complex words: For the positive prompts, we employ the set {satisfactory, excellent, perfect}, and for the negative prompts, the set {terrible, unsatisfactory, subpar} is utilized. The outcomes are depicted in Table XI. From this table, we observe that using complex prompts yields performance results comparable to those achieved with simple prompts. This suggests that the contributions of complex and simple prompts within the prompt-ensemble strategy are not significantly different.
| Dimensions | Single | Pair | Assessment | ||
[…]
| Without Q-Instruct | |||||
[…]
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 60.07% | 3.21 | 52.25% | 3.22 | 0.424/0.445 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 61.40% | 3.47 | 52.05% | 3.44 | 0.444/0.474 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 64.35% | 4.21 | 51.11% | 3.51 | 0.541/0.581 |
| With Q-Instruct | |||||
[…]
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 67.42% | 3.82 | 53.17% | 3.41 | 0.691/0.722 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 70.43% | 4.00 | 54.44% | 3.57 | 0.649/0.677 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 70.37% | 4.25 | 53.21% | 3.66 | 0.686/0.711 |
[…]
| Best Close-Source Performance | 74.10% | - | 78.07% | - | -/- |
V-H4 IQA evaluation strategy for CLIP-ViT-Large-14
In Tab. VI, we compare the IQA performance of MLLMs with CLIP-ViT-Large-14, the visual backbone of the majority of MLLMs. Attempting to understand whether the new language part (LLM) can do better than the original language part of CLIP, we try to compare between CLIP and MLLMs in a relatively aligned setting. Firstly, noticing that most MLLMs will resize images into as their input sizes, we align this setting on CLIP, and ignore the strategies as proposed by [clipiqa]. Secondly, same as the strategy on MLLMs, we also apply softmax pooling between good and poor, as in the CLIP’s zero-shot classification format: a photo of good quality and a photo of poor quality. Besides the two alignments, the quality scores of CLIP-ViT-Large-14 are obtained as follows:
where represents the cosine similarity calculation function.
V-H5 Special IQA settings for Flan-T5-based InstructBLIP
For InstructBLIP [iblip] (Flan-T5-XL), different from the majority of LLaMA-based (or MPT-based Otter-v1) MLLMs, the two top-frequency tokens are high (89%) and low (8%) instead of the common goodpoor. Henceforth, based on our motivation to only modify the argmax into softmax and follow the default top-frequency output tokens of MLLMs, we replace the probabilities of goodpoor into those of highlow in Eq. 1 for T5, defined as follows:
As validated in our experiments (Table XIII, the highlow pair generally predicts better than goodpoor on majority of databases. The better performance on MLLM-specific top-frequency tokens by side validates the effectiveness of our methodology for MLLMs on IQA.
V-H6 Special validation protocol for CGIQA-6K
The CGIQA-6K [zhang2023subjective] dataset contains two separate sub-sets which consist of 3,000 game images and 3,000 movie images respectively, with different instructions for human annotators during its subjective experiments. Therefore, we validate the MLLMs’ assessment performance on the two sub-sets individually and average the results for the final exhibition. The results of NIQE and CLIP-ViT-Large-14 are also obtained under the same protocol for a fair comparison.
V-I MLLMs Finetuned with Low-level Dataset
Additionally, we have proceeded to conduct the corresponding experiments separately to compare the performance of MLLMs that are fine-tuned on low-level multi-modal dataset. Q-Instruct [q-instruct] is the first large-scale, low-level multi-modal dataset available, prompting our decision to fine-tune four popular MLLMs (LLaVA-v1.5 (Vicuna-v1.5-7B), LLaVA-v1.5 (Vicuna-v1.5-13B), mPLUG-Owl2 (LLaMA-7B), and InternLM-XComposer-VL (InternLM).) using this dataset. The experimental results are presented in Table XIV. From the table, we observe that fine-tuning with the low-level multi-modal dataset significantly enhances performance on single image tasks (all finetuned MLLMs are superior to the best open-source performance), but only slightly improves performance on image pair tasks. This suggests that while the low-level multi-modal dataset helps MLLMs acquire better low-level knowledge, it does not sufficiently enhance their comparison abilities, resulting in limited performance gains for image pair tasks.
V-J Statement on Data Contamination
The Q-bench+ contains three tasks, where the first two tasks, (A1) perception and (A2) description, are evaluated with our own datasets proposed with the paper. For these two tasks, the questions, answers, or low-level descriptions in the two datasets are not seen by any existing MLLMs. Half of LLVisionQA+ (i.e. the test subset) and full of LLDescribe+ labels are kept private, to avoid being added to the training sets of any MLLMs. We hope that this measure will allow Q-bench+ to have long-term significance as an indicator of low-level visual abilities.
For the third task, (A3) assessment, the situation is a bit more complicated. For open-source models as tested, almost all of them have provided their technical reports, and as far as we know, no image quality assessment (IQA) dataset has participated in the multi-modality training stages of them. While text knowledge about image quality assessment should have been injected to them (e.g. a blurry image is a low quality image) during their pure-language training stages, we think this should not be regarded as data contamination for IQA, because the images cannot be seen by a language model. Instead, they are important knowledge for MLLMs to better link particular visual attributes (blur) to human opinions (quality), which motivates us to explore MLLMs for these tasks.